Determine the midpoint of the segment whose endpoints are A(7, 3, -4) and B(-3, 1, 0).
(5, 1, -2)
(2, 0, 2)
(2, 2, -2)
(-5, -1, -2)

Answer by Alan3354(69443)

Determine the midpoint of the segment whose endpoints are
A(7,3,-4) and B(-3,1,0).
--------------------
Find the average of each axis
(7-3)/2 = 2
(3+1)/2 = 2
(-4+0)/2 = -2
--> (2,2,-2)
-----------
Don't put spaces between the numbers.